A dipole is centered at the origin of a coordinate system, and a small charged sphere is some distance away along the perpendicular bisector of the dipole. The particle carries a uniformly distributed charge of -3.0 nC, and experiences a 200 nN electric force in the positive y direction.

(a) If the dipole charge is 10 nC and the dipole separation is 20 mm, how far away from the dipole is the sphere?

(b) How is the dipole moment oriented?


1
Expert's answer
2021-09-03T08:25:01-0400

Find the distance from the dipole ends to the sphere:


"D^2=a^2+x^2"

The force between the green and blue charge and the sphere:


"F_1=k\frac{qQ}{D^2}=k\frac{qQ}{a^2+x^2}.\\\space\\
F_2=k\frac{qQ}{D^2}=k\frac{qQ}{a^2+x^2}."

The sine of the angle between the x-axis and the forces is


"\sin\theta=\frac{a}{D}=\frac{a}{\sqrt{x^2+a^2}}."

The resultant of 200 nN is


"R=F_1\sin\theta+F_2\sin\theta=\\\space\\
=\frac{2kqQ}{a^2+x^2}·\frac{a}{\sqrt{x^2+a^2}}=\frac{2kaqQ}{(a^2+x^2)^{\frac32}}."

Solve this equation using any method you prefer (I used Wolframalpha.com), the distance is


"a=0.3\text{ m}."

As we see, the moment is oriented clockwise.